# Demystifying Language Model Forgetting with Low-rank Example Associations

### v6

@@ 4.3 Results of Predicting and Mitigating Forgetting
Figure 5: Log perplexity ($\downarrow$) or Token F1 ($\uparrow$) over upstream data by replay example selection strategies. The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting). The dash lines show the results achieved by upweighting upstream examples according to their ground truth forgetting. * and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests on all fine-tuning tasks.
@@
Mitigating forgetting with the predicted forgetting. We leverage the online or offline predicted forgetting by the matrix completion algorithms to reweigh examples during replay following the procedure introduced in Sec. 4.2. Figure 5 summarizes log perplexity or Token F1 over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR-T does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples, shown in dash lines in Figure 5.
Table 4: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$. Costs that are minor are displayed in smaller fonts.
@@
Further discussions. We show in Appendix F that targeted replay also slightly improves new task learning perplexity. We evaluate fine-tuned OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix F. Although we observe slightly improved performance of replaying predicted examples to random or no replay on most forgotten tasks, we do not see statistical significance. Future works can mitigate downstream task forgetting with predicted forgetting with alternative algorithms that leverage past examples (Aljundi et al., 2017; Buzzega et al., 2020). Appendix H demonstrates mild change in the pattern of forgetting in the held-out upstream examples when replay is applied. Our results in Appendix I show that matrix completion and embedding based methods can be combined to better improve forgetting prediction performance.
@@ 6 Conclusions
In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning. We showed that simple low rank patterns are dominant in the example associations and compared the complexity of the associations across model types and sizes. We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks. We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweigh examples for replay with predicted forgetting. Future works can extend the study to a continual learning setup where new domains or tasks are learned sequentially. We discuss limitations of the work in Appendix A.
@@ Appendix A Limitations
Although our work tries to be extensive, we consider there are limitations that can be taken as future works.
@@
Theoretical analysis of low-rank example associations. While we conclude that low-rank example associations are prevalent in LLM forgetting from empirical results, we do not have a theoretical guarantee on why example associations must be low rank, and in which cases the example associations become more complicated. We leave a theoretical understanding of low-rank example associations as future work.
@@
Scope of the experiment setups. Our analysis is restricted to supervised fine-tuning of LMs; other learning setups such as reinforcement learning is not studied. Besides, although we try to cover a broad category of newly learned tasks, the coverage may not be extensive. Due to limitations in computational resources, it was also not possible to repeat experiments under extensive hyperparameter configurations (e.g., learning rate schedulers), which we leave for future work. Nevertheless, we provide an analysis of more learning rate and batch size setups in Appendix H.
@@
Requirement of open access to upstream data. Our analysis requires open access to upstream data to evaluate their forgetting. As a result, our analysis is limited to open-data LLMs. In addition, similar to almost every replay-based algorithm, our forgetting mitigation algorithm requires access to upstream data. We hope that our work can inspire researchers that build closed-data LLMs with in-house data.
@@
Scale of the upstream examples. Due to computational resource restrictions, we significantly subsampled upstream pretraining corpora of LLMs in our analysis. Although expanding the set of upstream examples would not theoretically affect the low-rank observations, we decide to leave larger scale analysis as future works. Besides, we consider that matrix completion based forgetting prediction permits sparse association matrices, where upstream example forgetting $z_{ij}$ are sparsely collected, which we will explore in future work.
@@
Separating out intended forgetting from unintended forgetting. Aligning with most prior works on continual learning, our experiments focus on mitigation of forgetting. We consider the goal aligns with our experiment setup, as upstream examples from Dolma and Tulu are carefully curated and cleaned, and the newly learned tasks intend to accumulate knowledge instead of overwrite previous knowledge. However, we highlight that forgetting is not always detrimental; in fact, the models are often expected to forget (unlearn) sensitive, noisy, and outdated examples (Nguyen et al., 2022). We believe that our analysis of forgetting helpful for future works that performs targeted forgetting of knowledge.
@@ Appendix C Dataset, Model, and LM Training Details
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 12 and 13. We also annotate tasks used as in-domain training or test tasks.
Training and evaluation details. For full-parameter fine-tuning of non-instruction-tuned LLMs of all types, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$. The learning rate is chosen among {$1e^{-6}$, $2e^{-6}$, $5e^{-6}$} that achieves best average validation perplexity after fine-tuning OLMo-7B on 5 randomly chosen tasks from FLAN. For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA and Dolly, considering the small size of the training sets, we train the models only for 100 steps with an effective batch size of 8. For OLMo2-SFT-Mix tasks, we train the model for 1,000 steps. We use HuggingFace Transformers library for training and VLLM library for efficient inference. Due to the computational resource limitations, the statistics of forgetting are collected in a single run.
@@
Computational Infrastructure. We used 4 Quadro RTX A6000 GPUs for fine-tuning LLMs, and used 1 Quadrio RTX A6000 GPU for LLM inference.
@@
(c) Negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$)
Figure 7: A side-by-side comparison between the matrices of forgetting, inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$), and the negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$) we examined in Sec. 3.3.
@@ Appendix D Details of Forgetting Prediction and Replay
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 12 and 13. The train-test split for the in-domain tasks is randomly generated.
Training and evaluation details. We use Surprise Library 1.1.3222https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, MF, and KNN prediction models. For MF, we set the dimension of the learnable features (rank) as 5. We train the regression models for 1,000 epochs over the association matrices.
For in-domain test splits, we randomly sample 30 upstream examples and assume ground truth forgetting is known for these examples. This is required for predicting forgetting on the rest of upstream examples by additive linear, MF, and KNN methods. We repeat the experiment 10 times.
We used all-distilroberta-v1 sentence transformer model as the encoder in the implementation of the embedding similarity based prediction model of forgetting. At inference, given an upstream example $x_{j}$, we compute the averaged dot-product with all examples in the learned task $T_{i}$. We note that at inference time the approach does not require ground truth forgetting of a small number of examples. For a fair comparison with other matrix completion methods, we replace the prediction of the approach with ground truth forgetting on these examples.
@@
Figure 9: Reconstruction of $Z$ in OLMo-7B experiments with $k$-th singular value and vectors. Components of higher values of $k$ capture finer-grained details in $Z$.
@@
Figure 10: Reconstruction of log-perplexity measured forgetting association matrix $Z$ in OLMo-7B-Instruct experiments with $k$-th singular value and vectors. Higher values of $k$ capture finer-grained details in $Z$.
@@ Appendix E Details of the Example Similarity Metrics
Textual similarity. We calculate the textual cosine similarity between learned tasks and upstream examples using TF-IDF vectorized representations of training examples in each learned task $T_{i}$ and an upstream example $x_{j}$. We also measure text representation similarity with final layer representations of OLMo-1B.
@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
Table 6: Validation log perplexity of the learned tasks after fine-tuning LMs with different replay strategies.
@@
LLM leaderboard task evaluation. We evaluate downstream task performance of OLMo-7B-Instruct after fine-tuning on 4 held-out tasks from OLMo2-SFT-Mix (marked in Table 7). We evaluate on Open LLM Leaderboard tasks [^3] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about. For fine-tuned models, we compare no replay, replaying random examples, and replaying with forgetting predicted by offline MF. Tables 7 summarize the results.
We notice that fine-tuning OLMo-7B-Instruct over new tasks without any replay improves metrics scores on some tasks (MUSR, GPQA) but causes forgetting on other tasks (MMLU, BBH, IFEval). Among tasks where performance improved, we do not see benefits of example replay. Among tasks with decreased performance, we see offline-MF mitigates forgetting compared to no replay or random replay slightly.
We conjecture that replay-based approaches are not sufficient to significantly mitigate forgetting on their own, and can be combined with other approaches. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
@@
Table 8: Semantic meaning of $k$-th component in the factorization of the example association matrix $Z$. We identify top relevant learned tasks and upstream example domains to $k$-th component in the factorization of the example association matrix $Z$.

### v7

@@ 4.3 Results of Predicting and Mitigating Forgetting
Figure 5: Log perplexity ($\downarrow$) or Token F1 ($\uparrow$) over upstream data by replay example selection strategies. The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting). The dashed lines show the results achieved by upweighting upstream examples according to their actual forgetting after fine-tuning without replay. * and ** indicate significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests on all fine-tuning tasks.
@@
Table 4: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs MCMC. Costs that are minor are displayed in smaller fonts.
@@
Mitigating forgetting with the predicted forgetting. We leverage the online or offline predicted forgetting by the matrix completion algorithms to reweigh examples during replay following the procedure introduced in Sec. 4.2. Figure 5 summarizes log perplexity or Token F1 over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that example importance defined in these works are different from how easily the examples are forgotten. We also notice that MIR-T does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples, shown in dash lines in Figure 5.
@@
Effect of replay new task learning and patterns of forgetting. We show in Appendix F that targeted replay also slightly improves new task learning perplexity. Appendix H demonstrates mild change in the pattern of forgetting in the held-out upstream examples when replay is applied.
Further discussions. We evaluate fine-tuned OLMo-7B and OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix F and visualize the corresponding forgetting association matrix in Fig. 11. Although we observe slightly improved performance of replaying predicted examples to random or no replay on most forgotten tasks, we do not see statistical significance. Future works can mitigate downstream task forgetting with predicted forgetting with alternative algorithms that leverage past examples (Aljundi et al., 2017; Buzzega et al., 2020). Our results in Appendix I show that matrix completion and embedding based methods can be combined to better improve forgetting prediction performance. In Appendix J, we examine patterns of forgetting by fine-tuning over noisy or adversarial training data.
@@ 6 Conclusions
In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning. We showed that simple low rank patterns are dominant in the example associations and compared the complexity of the associations across model types and sizes. We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks. We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweight examples for replay with predicted forgetting. We discuss limitations and future work in Appendix A.
@@ Appendix A Limitations
Although our work aims to be extensive, we consider there are several limitations that can be taken as future work.
@@
The scope of the definition of forgetting. We mainly measure forgetting in log perplexity increase or exact match over upstream training data of LLMs. The choice of the metrics is based on the considerations that (1) they are connected to theoretical approximations of forgetting with first-order techniques (Sec. 3.3), allowing proper comparison, and (2) we aim to analyze forgetting on upstream examples that the models have seen before and minimize confounders such as generalization effects on unseen data.
Nevertheless, we note that the forgetting can be defined in many possible ways, each with different focus. Some measures of forgetting, such as an aggregated number of performance drop on domain-specific downstream tasks, are often more sensible to downstream LLM users. Some existing works also separate out “shallow forgetting” of language models, where the forgetting is easily recovered by twisting the prompts (Kotha et al., 2024; Chen et al., 2024). Importantly, forgetting defined in different ways may not always align with each other. We again remind the readers of the definition of forgetting used in the paper, and we consider analysis under different definitions of forgetting to be meaningful future work.
@@
Theoretical analysis of low-rank example associations. While we conclude that low-rank example associations are prevalent in LLM forgetting from empirical results, we do not have a theoretical guarantee on why example associations must be low rank, and in which cases the example associations become more complicated. We leave a theoretical understanding of low-rank example associations as future work. Besides, we consider that the line of research on mechanistic interpretability of LLMs (Conmy et al., 2023), which dissects subsets of neurons or computation graphs that are crucial for a prediction, can offer meaningful insights for interpreting the low-rank property.
@@
Scope of the experiment setups. Our analysis is restricted to supervised fine-tuning of LMs; other learning setups such as reinforcement learning are not studied. Besides, although we try to cover a broad category of newly learned tasks, the coverage may not be extensive. Due to limited computational resources, it was also not possible to repeat experiments under extensive hyperparameter configurations (e.g., learning rate schedulers), which we leave for future work. Nevertheless, we provide analysis under more learning rate and batch size setups in Appendix H.
@@
Requirement of open access to upstream data. Our analysis requires open access to upstream data to evaluate forgetting over them. As a result, replication of our analysis can be limited to LLMs trained on open data or in-house data that can be accessed. In addition, similar to almost every replay-based algorithm, our forgetting mitigation algorithm requires access to upstream data.
@@
Scale of the upstream examples. Due to computational resource restrictions, we significantly subsampled upstream pretraining corpora of LLMs in our analysis. Although expanding the set of upstream examples would not theoretically affect the low-rank observations, we decided to leave larger scale analysis as future work. Besides, we consider that matrix completion based forgetting prediction permits sparse association matrices, where upstream example forgetting $z_{ij}$ is sparsely collected, which is a feasible path towards generalizing the forgetting prediction framework to larger scale setups.
@@
Separating out intended forgetting from unintended forgetting. Aligning with most prior works on continual learning, our experiments focus on the mitigation of forgetting. We consider the goal aligns with our experiment setup, as upstream examples from Dolma and Tulu are carefully curated and cleaned, and the newly learned tasks intend to accumulate knowledge instead of overwrite previous knowledge. However, we highlight that forgetting is not always detrimental; in fact, the models are often expected to forget (unlearn) sensitive, noisy, and outdated examples (Nguyen et al., 2022; Blanco-Justicia et al., 2025). In addition, active forgetting can serve as a strategy to enhance the plasticity of models when learning new knowledge (Chen et al., 2023). We believe that our analysis of forgetting is helpful for future work that performs targeted forgetting of knowledge.
@@ Appendix C Dataset, Model, and LM Training Details
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 15 and 16. We also annotate tasks used as in-domain training or test tasks.
Training and evaluation details. For full-parameter fine-tuning of non-instruction-tuned LLMs of all types, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$. The learning rate is chosen among $1e^{-6}$, $2e^{-6}$, $5e^{-6}$} that achieves the the best average validation perplexity after fine-tuning OLMo-7B on 5 randomly chosen tasks from FLAN. For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA and Dolly, considering the small size of the training sets, we train the models only for 100 steps with an effective batch size of 8. For OLMo2-SFT-Mix tasks, we train the model for 1,000 steps. We use HuggingFace Transformers library for training and VLLM library for efficient inference. Due to computational resource limitations, the statistics of forgetting are collected in a single run.
@@
Computational Infrastructure. We used 4 Quadro RTX A6000 GPUs for fine-tuning LLMs, and used 1 Quadro RTX A6000 GPU for LLM inference.
@@
(c) Negated inner products of gradients (-$z_{ij}^{\textrm{g-g}}$)
Figure 7: A side-by-side comparison between the matrices of forgetting, inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$), and the negated inner products of gradients (-$z_{ij}^{\textrm{g-g}}$) we examined in Sec. 3.3.
@@ Appendix D Details of Forgetting Prediction and Replay
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 15 and 16. The train-test split for the in-domain tasks is randomly generated.
Table 5: Mean and standard deviation of forgetting, and the percentage where the forgetting $z_{ij}$ is non-negative, aggregated over the matrix $Z$. The table uses the same experiment setup as Figure 3.
Models / Statistics Avg. Forgetting Std. Forgetting Non-Neg %
OLMo-1B 0.0261 0.0425 90.36
OLMo-7B 0.0804 0.0626 96.29
MPT-7B 0.0917 0.0870 98.32
OLMo-7B-Instruct 0.0637 0.3700 49.61
OLMo2-7B 0.0045 0.0086 81.77
OLMo2-13B 0.0038 0.0078 80.11
Pythia-1B 0.1248 0.1406 98.53
Pythia-3B 0.0425 0.0523 89.24
Pythia-6.9B 0.1408 0.1517 99.20
Pythia-12B 0.1555 0.1691 99.03
Training and evaluation details. For MF, we set the dimension of the learnable features (rank) as 5. We train the regression models for 1,000 epochs over the association matrices.
For in-domain test splits, we randomly sample 30 upstream examples and assume the ground truth forgetting is known for these examples. This is required for predicting forgetting on the rest of upstream examples by additive linear, MF, and KNN methods. We repeat the experiment 10 times.
For forgetting prediction based on SBERT embedding similarity, we used all-distilroberta-v1 as the encoder. We also experiment with text-embedding-3-small model by OpenAI. The only trainable part of the prediction models is the MLP layers on top of the text embeddings. We train the models for 2,000 epochs. We note that at inference time the approach does not require ground truth forgetting of a small number of examples.
@@
Figure 9: Decomposition of $Z$ in OLMo-7B experiments with $k$-th singular value and vectors. Components of higher values of $k$ capture finer-grained details in $Z$.
@@
Figure 10: Decomposition of log-perplexity measured forgetting association matrix $Z$ in OLMo-7B-Instruct experiments with $k$-th singular value and vectors. Higher values of $k$ capture finer-grained details in $Z$.
@@ Appendix E Details of the Example Similarity Metrics
Token similarity. We calculate the textual cosine similarity between the learned tasks and upstream examples using TF-IDF vectorized representations of training examples in each learned task $T_{i}$ and an upstream example $x_{j}$.
Representational similarity. We measure textual representation cosine similarity given by the (1) final layer representations of OLMo-1B, (2) all-distilroberta-v1 sentence transformer (SBERT) model (Reimers & Gurevych, 2019), and (3) OpenAI text-embedding-3-small model. We average representations of maximum 100 training examples as the representation of a task $T_{i}$.
@@
(a) OLMo-7B, full association matrix of forgetting on LLM leaderboard tasks
(b) OLMo-7B, forgetting on LLM leaderboard tasks, $Z_{1}$
(c) OLMo-7B, forgetting on LLM leaderboard tasks, $Z_{2}-Z_{1}$
(d) OLMo-7B, forgetting on LLM leaderboard tasks, $Z_{3}-Z_{2}$
(e) OLMo-7B, forgetting on LLM leaderboard tasks, $Z_{4}-Z_{3}$
Figure 11: The forgetting association matrix $Z$ between $M=85$ learned tasks and $N=1,000$ examples from the LLM leaderboard tasks. The forgetting is defined as a prediction turning from correct to incorrect after fine-tuning. We further show the differences $Z_{r}-Z_{r-1}$ between rank-$r$ and rank-r1r-1 approximations. We note that $Z$ is binary, and we define the rank $r$ approximation $Z^{r}$ as $\sigma(\sum_{k=1}^{r}\bm{\alpha}_{k}\bm{\beta}_{k}^{T})$, where $\sigma$ is the sigmoid function.
@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
Table 7: Validation log perplexity of the learned tasks after fine-tuning LMs with different replay strategies. *: $p<0.05$ improvement compared to no replay.
@@
LLM leaderboard task evaluation. We evaluate the change of performance on Open LLM Leaderboard tasks [^1] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about before and after fine-tuning a model. We use same set of LLM leaderboard tasks as the OLMo technical report (Groeneveld et al., 2024). For OLMo-7B models, we evaluate downstream performance of 8 models fine-tuned the Dolly tasks. For OLMo-7B Instruct, We evaluate downstream performance after fine-tuning on 4 held-out tasks from OLMo2-SFT-Mix (marked in Table 9). We compare LLM leaderboard task performance of no replay, replaying random examples, and replaying with forgetting predicted by offline MF. Table 8 summarizes results on OLMo-7B; Table 9 summarizes results on OLMo-7B-Instruct.
We notice that fine-tuning OLMo-7B on Dolly improves downstream performance on most of the downstream tasks. This aligns well with the purpose of fine-tuning a LM that is not instruction-tuned. Nevertheless, we notice performance degradation on two of the tasks, namely ARC-Easy and Sciq, which indicates forgetting. Although offline MF Replay achieves higher accuracy scores on these two tasks compared to no-replay, we do not find the improvement statistically significant.
We see fine-tuning OLMo-7B-Instruct over new tasks without any replay improves metrics scores on some tasks (MUSR, GPQA) but causes forgetting on other tasks (MMLU, BBH, IFEval). Among tasks where performance improved, we do not see benefits of example replay. Among tasks with decreased performance, we see offline-MF mitigates forgetting compared to no replay or random replay only very marginally.
We conjecture that replay-based approaches are not sufficient to significantly mitigate forgetting on their own, and can be combined with other approaches. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work.
Table 8: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
MF-Offline 67.52 42.35 74.50 77.56 44.25 79.97 85.10 67.17
@@
Appendix G Towards Interpreting Fine-Grained Associations
Table 10: Semantic meaning of the $k$-th component in the factorization of the example association matrix $Z$. We identify top relevant learned tasks and upstream example domains to the $k$-th component in the factorization of the example association matrix $Z$.